# Higher-order triadic percolation on random hypergraphs

### v1

@@ III.2 Higher-order triadic percolation as a fully-fledged dynamical process
Higher-order triadic percolation defines a fully-fledged dynamical process in which the fraction of node in the giant component $R^{(t)}$ is in general a non-trivial function of time. In order to see this we can encode the dynamical Eqs. 5-7 implementing Step 1 and determining $R^{(t)}$ once $p_{H}^{(t-1)}$ is known as a function
@@
Note that the function $f_{m}(p_{H}^{(t-1)})$ already encodes for the result of the percolation process, thus encodes already for the percolation phase transition of hyperedge percolation. On the other side Step 2 is determined by the Eq. 8 that evidently can be expressed as a function determining $p_{H}^{(t)}$ once $R^{(t)}$ is known, i.e.
@@
This one-dimensional map, defines the evolution of the percolation order parameter $R$ of HOTP as a function of time, i.e. $R=R^{(t)}$ for an infinite random hypergraph with HOTI. This one dimensional map thus encode the dynamical nature of HOTP, and can be used to predict the dynamical behavior of the order parameter. Indeed, by iterating the map we can build a Cobweb plot that accurately predicts the dynamics of the order parameter on large hypergraphs (see Fig. 2). As we will discuss below and prove in Appendix A, this map undergoes a route to chaos in the universality class of the logistic map. In particular the dynamics of the order parameter can either go to a fixed point $R=R^{\star}$, corresponding to a static asymptotic state of HOTP, or it can display periodic oscillations or even display a chaotic dynamics in the thermodynamic limit (see Fig. 2).
The critical behavior of the dynamics can be studied by analysing its derivative $J$. A stable stationary solution $R^{\star}$ of Eq. 11 satisfy the fix point equation
@@
Thus the stationary solution loses its stability for $|J|=1$. When $J=1$, as long as $R=R^{\star}>0$, we observe a discontinuous hybrid transition with a square root singularity. When $J=-1$ we observe instead a bifurcation, resulting in period-2 oscillations. Similarly, the critical point of the onset of period-4 oscillations can be obtained by analysing the derivative $J_{2}$ of the second-iterate function $h^{2}_{m,p}=h_{m,p}\circ h_{m,p}$. The bifurcation of period-2 oscillations takes place at $J_{2}=-1$, resulting in period-4 oscillations.
In the next paragraphs we will discuss how this these theoretical insights shed light into the critical properties of HOTP.
@@ III.3 Critical properties of higher-order triadic percolation
In order to investigate the critical properties of HOTP we consider the effect induced by the sign of the HOTI. Thus we will discuss the scenarios in which only positive HOTIs are present, the case in which only negative HOTIs are present, and the case in which both positive and negative HOTIs are present with non vanishing probability.
III.3.1 Higher-order triadic percolation in absence of negative HOTIs
If only positive HOTIs are present, the Eq. 8 determining the regulatory step (Step 2) reduces to
@@ III.3.1 Higher-order triadic percolation in absence of negative HOTIs
In this case we cannot have a route to chaos because the map $h_{m,p}(R)=f_{m}(g_{p}(R))$ is monotonically increasing with $R$ [50] (see Fig. 3 for an example). Indeed we have
@@
This stationary state, $R^{\star}$ displays a discontinuous hybrid phase transition between a zero value $R^{\star}=0$ to a non zero value $R=R^{\star}>0$ at the critical point $p=p_{c}$.
@@
In order to show that the transition is discontinuous an hybrid, consider expand the stationary point equation around the critical point $p=p_{c}$. Let us indicate with $\delta p$ the distance from the critical point, i.e. $0<\delta p=p-p_{c}\ll 1$ and with $\delta R$ the corresponding change in the order parameter $0<\delta R=R^{\star}(p)-R^{\star}(p_{c})\ll 1$. Assuming that $h_{m,p}(R)$ is differentiable up to the second order at $R_{c}=R^{\star}(p_{c})$, we can expand Eq. 17 obtaining
$\delta R=h_{m,p_{c}}^{\prime}\left(R_{c}\right)\delta R+\frac{1}{2}h_{m,p_{c}}^{\prime\prime}\left(R_{c}\right)(\delta R)^{2}+\frac{\partial h_{m,p_{c}}\left(R_{c}\right)}{\partial p}\delta p.$
Since $h_{m,p_{c}}^{\prime}\left(R_{c}\right)=1$ at $p=p_{c}$, we obtain
$\frac{1}{2}h_{p_{c}}^{\prime\prime}\left(R_{c}\right)(\delta R)^{2}+\frac{\partial h_{p_{c}}\left(R_{c}\right)}{\partial p}\delta p=0.$ (19)
Thus since $h_{p_{c}}^{\prime\prime}\left(R_{c}\right)$ and ${\partial h_{p_{c}}\left(R_{c}\right)}/{\partial p}$ are finite and have opposite signs, the order parameter $R$ has a square-root singularity for $p\to p_{c}^{+}$, i.e.
@@
indicating that the transition is hybrid. Our theoretical predictions perfectly match our simulation results (see Fig. 4). In particular we have considered a random hypergraph with fixed hyperedge cardinality $m$ (i.e. with $Q(m^{\prime})=\delta_{m^{\prime},m}$ and Poisson distribution $P(k)$ and $\hat{P}(\kappa)$. The hypergraph with larger hyperedge cardinality $m$ is more robust and displays simultaneously a lower critical threshold $p_{c}$, and a smaller discontinuity $R^{\star}(p_{c})$ (see Fig. 4(b)). This result is in line with the result on simple hyperedge percolation [14], where hypergraphs with larger hyperedge cardinality $m$ are more robust, albeit in that case the percolation transition is continuous.
Figure 4: Phase diagram of random hyperedge percolation in absence of negative HOTIs. Panel (a) shows the size of the giant component $R$ as a function of $p_{H}$. Solid lines indicate theoretical predictions (Eq. 6 - Eq. 8) symbols indicate the results of Monte Carlo simulations. Panel (b) shows the numerical calculation of critical threshold $p_{c}$ and the non-trivial fixed point $R^{\star}(p_{c})$ at $p=p_{c}$ as a function of the hyperedge cardinality $m$. The Monte Carlo simulations are performed on hypergraphs with HOTI having $N=10^{5}$ nodes. The structural hypergraph has hyperedges of fixed cardinality $m$ and has Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=4$. The degree distribution $\hat{P}^{+}(\kappa)$ of is Poisson with an average positive regulatory degree $c^{+}=2$.
III.3.2 Higher-order triadic percolation in absence of positive HOTIs
In this paragraph, we consider the other limiting case in which HOTP includes only negative HOTIs. In this case a hyperedge is down-regulated if at least one of is negative regulator nodes is active. This scenario can be interpreted as HOTP in the limit in which the positive regulatory interactions a very large. Thus in Step 2 of HOTP we substitute Eq. 8 with
@@ III.3.2 Higher-order triadic percolation in absence of positive HOTIs
Indeed, according to our interpretation, this equation can be obtained from Eq. 8 by putting $G_{0}^{+}\left(1-R^{(t)}\right)$ identically equal to zero.
Also in this case, similarly to the precedent limiting case, the map $h_{m,p}(R)$ is monotonic, thus we do not observe a route to chaos (see Figure 3). However, it can be easily show that, differently from the precedent limiting case, in absence of any positive interactions the map $h_{m,p}(R)$ is monotonically decreasing. In fact if we have
@@
It follows that the only critical points of the dynamic equations can be a period-2 bifurcation for $J=-1$. This can occur at $R^{\star}=0$ or at $R^{\star}>0$. Note however that we can additionally also observe continuous phase transitions from a stationary solution $R^{\star}=0$ to a stationary solution $R^{\star}>0$. This transition will not occur at a specific value of $J$, and can occur as long as $|J|<1$, condition that ensures the stability of the stationary solution. Indeed the continuous transition occurs whereas the hyperedge percolation equation displays a phase transition, as the change of solutions between $R^{\star}=0$ and $R^{\star}>0$ is already encoded in the function $f_{m}(p_{H})$. This happens when the maximum of $p_{H}=g_{p}(R)$ which is reached at $R=0$ equals to the critical threshold of hyperedge percolation $p_{H}^{c}$ provided that $|J|<1$ (see Fig. 5.(a)). Interestingly, in this scenario, the threshold for the continuous transitions is independent of the regulatory network and only depends on the structural hypergraph.
@@
By studying the the iteration of the map $h_{m,p}\circ h_{m,p}(R)$ it can be shown that period-4 bifurcations are not possible. Thus the characterization of the critical properties of HOTP in absence of positive regulation reveals that in this case only period-2 bifuractions and continuous transitions can be observed.
In Fig 6 we compare our theoretical predictions with Monte Carlo simulations performed on hypergraphs with fixed cardinality $m$ and Poisson hyperdegree distribution $P(k)$. From this figure we can observe the presence of two period-2 bifurcations of small value of $m$ which at the tricritical point disappear giving rise to a continuous phase transition, as predicted by our theory (see Fig. 6 (d)). Moreover, also in this case, as in the scenario where only positive regulation are present, we observe that for hypergraph with hyperedges of fixed cardinality $m$, hypergraphs with larger value of $m$ are more robust. In fact, for larger values of $m$, the onset of period-2 oscillation occurs at smaller values of $p$, and collapse of the network (onset of period-2 oscillations or continuous phase transition at $R=0$) also occurs at a values of $p$ that decreases with $m$.
@@ III.3.3 Route to chaos in higher-order triadic percolation
When both positive and negative HOTIs are present, the order parameter of HOTP undergoes a route to chaos in the universality class of the logistic map and HOTP becomes a fully-fledged dynamical process (see Fig. 7). This can be demonstrated by considering the map $R^{(t)}=h_{m,p}(R^{(t-1)})$ and, using the results of Feigenbaumm [49, 50], showing that the map is continuous, unimodal and displays a quadratic maximum conditions that define the logistic map universality class of the route to chaos (see Fig. 3 for an example and Appendix A For a detailed discussion). This phenomenology is in line to what happens for standard triadic percolation [13].
@@
Interestingly, the increase of random deactivation of hyperedges (i.e. the decrease of $p$) does not result in a monotonic suppression of the giant component. Instead, the order parameter $R$ in some cases displays a non-monotonous behavior as a function of $p$ and reentrant phase transitions (see Fig. 8). This consist on a collapsed state with $R^{(t)}=0$ for all values of $t$, observed in a ranges of values of $p\in[p_{-},p_{+}]$, while for $p<p_{-}$ and for $p>p_{+}$ the giant component is non zero. The reentrant nature of the phase implies the naïvely not intuitive result that for a higher level of random damage (smaller values of $p$) the giant component can be restored. Reentrant phase transitions are also observed in correlated multilayer networks [48] and in multilayer networks with extended-range percolation [51], however in all these example the fraction of nodes in the giant component is unique in the thermodynamics limit, while this is an example of a reentrant orbit diagram of percolation. Finally we note that such phenomenon can be also observed on networks, (i.e. for $Q(m)=\delta_{m,2}$), under suitable conditions.
@@
Figure 7: In presence of both positive and negative HOTIs, the phase diagram of HOTP reduces to an orbit diagram displaying a route to chaos in the universality class of the logistic map. Panel (a) displays the theoretical prediction of this orbit diagram and panel (b) displays the results of Monte Carlo simulations. The structural hypergraph has $N=10^{6}$ nodes, hyperedges of fixed cardinality $m=10$ and Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=10$. The Poisson degree distribution $\hat{P}^{\pm}(\kappa)$ with an average $c^{+}=3$ and $c^{-}=5.5$.
Figure 8: We provide two examples of orbit diagrams showing the reentrant phase transition on HOTP, which the hypergraph can be re-activated by decreasing the value of $p$, i.e. increasing the probability of random deactivation of the hyperedges. Panels (a) and (b) show two examples of theoretical orbit diagrams of HOTP displaying the reentrant phase transition. Panels (c) and (d) show the orbit diagrams obtained by implementing the corresponding Monte Carlo simulations. Panels (e) and (f) reveals that Eq. 30, i.e. the intersection between the curve $y=h_{m,p}(h_{m,p}(\tilde{R}))-R^{\star}$ and the orange dash line indicates $y=0$, provides the correct prediction of the value of $p$ at which the either the hypergraph collapse of the reentrant phase transition are observed. In both examples, the structural hypergraph has $N=5\times 10^{5}$ nodes, Poisson hyperdegree distribution $P(k)$ with average $c=30$ and hyperedges with fixed cardinality given by $m=2$ (in panels (a),(c),(e)) and $m=3$ (in panels (b),(d),(f)). The Poisson degree distributions $\hat{P}^{\pm}(\kappa)$ have average positive and negative regulatory degree given by $c^{+}=1,c^{-}=2.9$ (in panels (a),(c),(e)); and $c^{+}=1,c^{-}=4$ (in panels (b),(d),(f)).
@@
Let us provide the theoretical understanding for the occurrence of the reentrant collapsed phase. Specifically, let us investigate the conditions for observing a non-trivial dynamics of the order parameter. In order to illustrate the argument, let us consider a continuous unimodal map $h_{m,p}(R)$ (see Fig. 9) with maximum $\tilde{R}$ having compact support on a proper subinterval of $[0,1]$. Furthermore, let us assume that the map has two fixed points of $R=h_{m,p}(R)$ denoted $\underline{R^{\star}}$ and $\overline{R^{\star}}$ with $\underline{R^{\star}}<\overline{R^{\star}}$ where $\underline{R^{\star}}$ is an unstable fixed point. Let us denote with $\overline{R}$ the value of $R$ with $R>\underline{R^{\star}}$ satisfying $h_{m,p}(\overline{R})=\underline{R^{\star}}$. This scenario is consistent for example (see Fig. 9) with the properties of the map $h_{m,p}(R)$ for a random hypegraph with Poisson hyperdegree distribution $P(k)$ with average hyperdegree $c$ and constant hyperedge cardinality $m$ where the regulatory degree distribution $\hat{P}(\kappa^{\pm})$ are Poisson distributions with averages $c$ and $c^{\pm}$ respectively. In this case the function $h_{m,p}(R)$ has a unique maximum at $R=\tilde{R}$ at
@@
Moreover $\left.\partial h/\partial R\right|_{R=\underline{R^{\star}}}>1$, thus $\underline{R^{\star}}$ is an unstable fixed point, while $\left.\partial h/\partial R\right|_{R=\overline{R^{\star}}}<1$, thus $\overline{R^{\star}}$ is a stable fixed point.
Under the stated conditions of the map $h_{m,p}$, if $R>\overline{R}$ or $R<\underline{R^{\star}}$, the dynamics will converge to the trivial collapsed state. In order to observe a non-trivial dynamics the interval $\left(\underline{R^{\star}},\overline{R}\right)$ must be mapped onto this itself under the action of the map $h_{m,p}$. Therefore, the condition for having non-trivial dynamics is given by

### v2

@@ III.2 Higher-order triadic percolation as a fully-fledged dynamical process
Higher-order triadic percolation defines a fully-fledged dynamical process in which the fraction of nodes in the giant component $R^{(t)}$ is in general a non-trivial function of time. In order to see this we can encode the dynamical Eqs. (5)-(7) implementing Step 1 and determining $R^{(t)}$ once $p_{H}^{(t-1)}$ is known as a function
@@
Note that the function $f_{m}(p_{H}^{(t-1)})$ already encodes for the result of the percolation process, thus encodes already for the percolation phase transition of hyperedge percolation. On the other side, Step 2 is determined by the Eq. (8) that evidently can be expressed as a function determining $p_{H}^{(t)}$ once $R^{(t)}$ is known, i.e.
@@
This one-dimensional map defines the evolution of the percolation order parameter $R$ of HOTP as a function of time, i.e. $R=R^{(t)}$ for an infinite random hypergraph with HOTI. This one-dimensional map thus encodes the dynamical nature of HOTP and can be used to predict the dynamical behavior of the order parameter. Indeed, by iterating the map we can build a Cobweb plot that accurately predicts the dynamics of the order parameter on large hypergraphs (see Fig. (2)). As we will discuss below and prove in Appendix A, this map undergoes a route to chaos in the universality class of the logistic map. In particular, the dynamics of the order parameter can either go to a fixed point $R=R^{\star}$, corresponding to a static asymptotic state of HOTP, or it can display periodic oscillations or even display a chaotic dynamics in the thermodynamic limit (see Fig. 2).
The critical behavior of the dynamics can be studied by analyzing its derivative $J$. A stable stationary solution $R^{\star}$ of Eq. (11) satisfy the fix point equation
@@
Thus the stationary solution loses its stability for $|J|=1$. When $J=1$, as long as $R=R^{\star}>0$, we observe a discontinuous hybrid transition with a square root singularity. When $J=-1$ we observe instead a bifurcation, resulting in period-2 oscillations. Similarly, the critical point of the onset of period-4 oscillations can be obtained by analyzing the derivative $J_{2}$ of the second-iterate function $h^{2}_{m,p}=h_{m,p}\circ h_{m,p}$. The bifurcation of period-2 oscillations takes place at $J_{2}=-1$, resulting in period-4 oscillations.
In the next paragraphs, we will discuss how these theoretical insights shed light on the critical properties of HOTP.
@@ III.3 Critical properties of higher-order triadic percolation
In order to investigate the critical properties of HOTP we consider the effect induced by the sign of the HOTI. Thus we will discuss the scenarios in which only positive HOTIs are present, the case in which only negative HOTIs are present, and the case in which both positive and negative HOTIs are present with non-vanishing probability.
III.3.1 Higher-order triadic percolation in the absence of negative HOTIs
If only positive HOTIs are present, the Eq. (8) determining the regulatory step (Step 2) reduces to
@@ III.3.1 Higher-order triadic percolation in the absence of negative HOTIs
In this case, we cannot have a route to chaos because the map $h_{m,p}(R)=f_{m}(g_{p}(R))$ is monotonically increasing with $R$ [52] (see Fig. 3 for an example). Indeed we have
@@
This stationary state, $R^{\star}$ displays a discontinuous hybrid phase transition between a zero value $R^{\star}=0$ to a non-zero value $R=R^{\star}>0$ at the critical point $p=p_{c}$.
@@
In order to show that the transition is discontinuous and hybrid [13, 39], consider expanding the stationary point equation around the critical point $p=p_{c}$. Let us indicate with $\delta p$ the distance from the critical point, i.e. $0<\delta p=p-p_{c}\ll 1$ and with $\delta R$ the corresponding variation of the order parameter $0<\delta R=R^{\star}(p)-R^{\star}(p_{c})\ll 1$. Using Eq.(17), and assuming that $h_{m,p}(R)$ is differentiable up to the second order at $R^{\star}(p_{c})$, we obtain
$\delta R=\frac{\partial h_{m}(R^{\star}(p_{c}),p_{c})}{\partial R}\delta R$ $+$ $\frac{1}{2}\frac{\partial^{2}h_{m}(R^{\star}(p_{c}),p_{c})}{\partial R^{2}}(\delta R)^{2}$
$+$ $\frac{\partial h_{m}\left(R^{\star}(p_{c}),p_{c}\right)}{\partial p}\delta p.$
Since Eq. (18) holds, $\partial h_{m}\left(R^{\star}(p_{c}),p_{c}\right)/\partial R=1$. Thus inserting this in the above equation we obtain
$\frac{1}{2}\frac{\partial^{2}h_{m}(R_{c},p_{c})}{\partial R^{2}}(\delta R)^{2}+\frac{\partial h_{m}\left(R_{c},p_{c}\right)}{\partial p}\delta p=0.$ (19)
Given that $\partial^{2}h_{m}(R^{\star}(p_{c}),p_{c})/\partial R^{2}$ and ${\partial h_{m}\left(R^{\star}(p_{c}),p_{c}\right)}/{\partial p}$ are finite and have opposite signs, the order parameter $R$ has a square-root singularity for $p\to p_{c}^{+}$, i.e.
@@
indicating that the transition is hybrid. Our theoretical predictions perfectly match our simulation results (see Fig. 4). In particular, we have considered a random hypergraph with fixed hyperedge cardinality $m$ (i.e. with $Q(m^{\prime})=\delta_{m^{\prime},m}$ and Poisson distribution $P(k)$ and $\hat{P}(\kappa)$. The hypergraph with larger hyperedge cardinality $m$ is more robust and displays simultaneously a lower critical threshold $p_{c}$, and a smaller discontinuity $R^{\star}(p_{c})$ (see Fig. 4(b)). This result is in line with the result on simple hyperedge percolation [14], where hypergraphs with larger hyperedge cardinality $m$ are more robust, albeit in that case the percolation transition is continuous.
Figure 4: Phase diagram of random hyperedge percolation in the absence of negative HOTIs. Panel (a) shows the size of the giant component $R$ as a function of $p_{H}$. Solid lines indicate theoretical predictions (Eq. (6) - Eq. (8)) symbols indicate the results of Monte Carlo simulations. Panel (b) shows the numerical calculation of critical threshold $p_{c}$ and the non-trivial fixed point $R^{\star}(p_{c})$ at $p=p_{c}$ as a function of the hyperedge cardinality $m$. The Monte Carlo simulations are performed on hypergraphs with HOTI having $N=10^{5}$ nodes. The structural hypergraph has hyperedges of fixed cardinality $m$ and has Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=4$. The degree distribution $\hat{P}^{+}(\kappa)$ of is Poisson with an average positive regulatory degree $c^{+}=2$.
III.3.2 Higher-order triadic percolation in the absence of positive HOTIs
In this paragraph, we consider the other limiting case in which HOTP includes only negative HOTIs. In this case, a hyperedge is down-regulated if at least one of is negative regulator nodes is active. This scenario can be interpreted as HOTP as the limit in which the positive regulatory degrees are infinite. Thus in Step 2 of HOTP we substitute Eq. (8) with
@@ III.3.2 Higher-order triadic percolation in the absence of positive HOTIs
Indeed, according to our interpretation, this equation can be obtained from Eq. (8) by putting $G_{0}^{+}\left(1-R^{(t)}\right)$ identically equal to zero.
Also in this case, similarly to the precedent limiting case, the map $h_{m,p}(R)$ is monotonic, thus we do not observe a route to chaos (see Figure 3). However, it can be easily shown that, differently from the precedent limiting case, in the absence of any positive interactions the map $h_{m,p}(R)$ is monotonically decreasing. In fact, if we have
@@
It follows that the only critical points of the dynamic equations can be a period-2 bifurcation for $J=-1$. This can occur at $R^{\star}=0$ or at $R^{\star}>0$. Note however that we can additionally also observe continuous phase transitions from a stationary solution $R^{\star}=0$ to a stationary solution $R^{\star}>0$. This transition will not occur at a specific value of $J$, but can occur as long as $|J|<1$, the condition that ensures the stability of the stationary solution. Indeed the continuous transition occurs whereas the hyperedge percolation equation displays a phase transition, as the change of solutions between $R^{\star}=0$ and $R^{\star}>0$ is already encoded in the function $f_{m}(p_{H})$. This happens when the maximum of $p_{H}=g_{p}(R)$, which is reached at $R=0$, equals the critical threshold of hyperedge percolation $p_{H}^{c}$ provided that $|J|<1$ (see Fig. 5.(a)). Interestingly, in this scenario, the threshold for the continuous transitions is independent of the regulatory network and only depends on the structural hypergraph.
@@
By studying the the iteration of the map $h_{m,p}\circ h_{m,p}(R)$ it can be shown that period-4 bifurcations are not possible. Thus the characterization of the critical properties of HOTP in the absence of positive regulation reveals that in this case only period-2 bifurcations and continuous transitions can be observed.
In Fig 6 we compare our theoretical predictions with Monte Carlo simulations performed on hypergraphs with fixed cardinality $m$ and Poisson hyperdegree distribution $P(k)$. From this figure, we can observe the presence of two period-2 bifurcations of small value of $m$ which at the tricritical point disappear giving rise to a continuous phase transition, as predicted by our theory (see Fig. 6 (d)). Moreover, also in this case, as in the scenario where only positive regulations are present, we observe that for hypergraphs with hyperedges of fixed cardinality $m$, hypergraphs with larger value of $m$ are more robust. In fact, for larger values of $m$, the onset of period-2 oscillation occurs at smaller values of $p$, and collapse of the network (onset of period-2 oscillations or continuous phase transition at $R=0$) also occurs at values of $p$ that decreases with $m$.
@@ III.3.3 Route to chaos in higher-order triadic percolation
When both positive and negative HOTIs are present, the order parameter of HOTP undergoes a route to chaos in the universality class of the logistic map, turning HOTP to a fully-fledged dynamical process (see Fig. 7). This can be demonstrated by considering the map $R^{(t)}=h_{m,p}(R^{(t-1)})$ and, using the results of Feigenbaum, [51, 52], which show that if the map is continuous, unimodal and displays a quadratic maximum, it will undergo the route to chaos in the universality class of the logistic map (see Fig. 3 for an example and Appendix A For a detailed discussion). This phenomenology is in line with what happens for standard triadic percolation [13].
Figure 7: In the presence of both positive and negative HOTIs, the phase diagram of HOTP reduces to an orbit diagram displaying a route to chaos in the universality class of the logistic map. Panel (a) displays the theoretical prediction of this orbit diagram and panel (b) displays the results of Monte Carlo simulations. The structural hypergraph has $N=10^{6}$ nodes, hyperedges of fixed cardinality $m=10$, and Poisson hyperdegree distribution $P(k)$ with an average hyperdegree $c=10$. The Poisson degree distribution $\hat{P}^{\pm}(\kappa)$ with an average $c^{+}=3$ and $c^{-}=5.5$.
Figure 8: The phase diagram of HOTP occurring on a real-world hypergraph displays a dynamical behavior that is qualitatively similar to the dynamical behavior of HOTP on random hypergraphs. The three panels display the phase diagram of HOTP in the presence of exclusively positive HOTI (panel (a)); in the presence of exclusively negative HOTI (panel(b)); in the presence of both positive and negative HOTIs (panel (c)). The structural hypergraph is the recipe ingredient hypergraph of $N=6714$ nodes and $E_{H}=39774$ hyperedges, with the largest hyperedge cardinality of 65. The data is freely available at the repository  [53]. The HOTIs are generated randomly between the nodes and the hyperedge of the real hypergraph and the average number $c^{+}$ and $c^{-}$ of positive and negative regulatory interactions is: $c^{+}=10,c^{-}=0$ (panel (a)); $c^{+}=\infty,c^{-}=15$ (panel (b)), and $c^{+}=10,c^{-}=15$ (panel (c)).
III.3.4 HOTP on real hypergraphs
An important question that arises is whether the observed critical and dynamical behavior of HOTP remains unchanged also on hypergraphs that are defined departing from the random hypergraphs that we have considered so far. Our results show that the predicted dynamical behavior of HOTP, derived analytically for random hypergraphs, is not restricted to random hypergraphs and can remain qualitatively unchanged also if HOTP is defined on real hypergraph architectures. In particular, to investigate the matter, we have considered a real-world hypergraph of recipe ingredients with $N=6714$ nodes representing ingredients and $E_{H}=39774$ hyperedges with the largest hyperedge cardinality of 65, representing recipes containing a combination of the ingredients, from the Repository [53]. On top of this real-world hypergraph, we have generated random regulatory networks determining the HOTIs between the nodes and the hyperedges. In Fig. 8 we show numerical evidence that HOTP occurring on the considered real hypergraph and studied in the presence of only positive, only negative, and in the presence of both positive and negative HOTIs reveals the same dynamical behavior of HOTP occurring on a random hypergraph. Our course does not exclude that other real-world hypergraphs might display a different behavior in particular if they are embedded into real space where we know that triadic percolation displays spatiotemporal dynamics of the giant component (see for instance discussion in Ref. [32]).
III.3.5 Reentrant orbit diagram in HOTP
@@
Interestingly, for some choices of the HOTp parameter values, the increase of random deactivation of hyperedges (i.e. the decrease of $p$) does not result in a monotonic suppression of the giant component. Instead, the order parameter $R$ in some cases displays a non-monotonous behavior as a function of $p$ and reentrant phase transitions (see Fig. 9). This consists of a collapsed state with $R^{(t)}=0$ for all values of $t$, observed in a range of values of $p\in[p_{-},p_{+}]$, while for $p<p_{-}$ and for $p>p_{+}$ the giant component is non zero. The reentrant nature of the phase implies the naïvely not intuitive result that for a higher level of random damage (smaller values of $p$) the giant component can be restored. Reentrant phase transitions are also observed in correlated multilayer networks [50] and in multilayer networks with extended-range percolation [54], however in all these example the fraction of nodes in the giant component is unique in the thermodynamics limit, while this is an example of a reentrant orbit diagram of percolation. Finally we note that such phenomenon can be also observed on networks, (i.e. for $Q(m)=\delta_{m,2}$), under suitable conditions.
@@
Figure 9: We provide two examples of orbit diagrams showing the reentrant phase transition on HOTP, in which the hypergraph can be re-activated by decreasing the value of $p$, i.e. increasing the probability of random deactivation of the hyperedges. Panels (a) and (b) show two examples of theoretical orbit diagrams of HOTP displaying the reentrant phase transition. Panels (c) and (d) show the orbit diagrams obtained by implementing the corresponding Monte Carlo simulations. Panels (e) and (f) reveals that Eq. (30), i.e. the intersection between the curve $y=h_{m,p}(h_{m,p}(\tilde{R}))-R^{\star}$ and the orange dash line indicates $y=0$, provides the correct prediction of the value of $p$ at which either the hypergraph collapse of the reentrant phase transition are observed. In both examples, the structural hypergraph has $N=5\times 10^{5}$ nodes, Poisson hyperdegree distribution $P(k)$ with average $c=30$ and hyperedges with fixed cardinality given by $m=2$ (in panels (a),(c),(e)) and $m=3$ (in panels (b),(d),(f)). The Poisson degree distributions $\hat{P}^{\pm}(\kappa)$ have average positive and negative regulatory degree given by $c^{+}=1,c^{-}=2.9$ (in panels (a),(c),(e)); and $c^{+}=1,c^{-}=4$ (in panels (b),(d),(f)).
@@
Let us provide the theoretical understanding of the occurrence of the reentrant collapsed phase. Specifically, let us investigate the conditions for observing non-trivial dynamics of the order parameter. In order to illustrate the argument, let us consider a continuous unimodal map $h_{m,p}(R)$ (see Fig. 10) with maximum $\tilde{R}$ having compact support on a proper subinterval of $[0,1]$. Furthermore, let us assume that the map has two fixed points of $R=h_{m,p}(R)$ denoted $\underline{R^{\star}}$ and $\overline{R^{\star}}$ with $\underline{R^{\star}}<\overline{R^{\star}}$ where $\underline{R^{\star}}$ is an unstable fixed point. Let us denote with $\overline{R}$ the value of $R$ with $R>\underline{R^{\star}}$ satisfying $h_{m,p}(\overline{R})=\underline{R^{\star}}$. This scenario is consistent for example (see Fig. 10) with the properties of the map $h_{m,p}(R)$ for a random hypergraph with Poisson hyperdegree distribution $P(k)$ with average hyperdegree $c$ and constant hyperedge cardinality $m$ where the regulatory degree distribution $\hat{P}(\kappa^{\pm})$ are Poisson distributions with averages $c$ and $c^{\pm}$ respectively. In this case the function $h_{m,p}(R)$ has a unique maximum at $R=\tilde{R}$ at
@@
Moreover, $\left.\partial h/\partial R\right|_{R=\underline{R^{\star}}}>1$, thus $\underline{R^{\star}}$ is an unstable fixed point, while $\left.\partial h/\partial R\right|_{R=\overline{R^{\star}}}<1$, thus $\overline{R^{\star}}$ is a stable fixed point.
Under the stated conditions of the map $h_{m,p}$, if $R>\overline{R}$ or $R<\underline{R^{\star}}$, the dynamics will converge to the trivial collapsed state. In order to observe a non-trivial dynamics, the interval $\left(\underline{R^{\star}},\overline{R}\right)$ must be mapped onto this itself under the action of the map $h_{m,p}$. Therefore, the condition for having non-trivial dynamics is given by